A reckless hunter is on a boat which is stationary in the middle of a lake.




They fire their gun horizontally 10 times. Each time they fire, a 12 g bullet leaves the gun at a speed of 410 m s−1.




What is the speed of the boat after they have have finished firing their gun if the total mass of the boat, hunter, and gun is 150 kg? (in m s−1 to 2 s.f)




Assume no drag forces from the water.





1
Expert's answer
2022-04-06T10:13:05-0400

Answer

the speed of the boat after they have have finished firing 

Is

"V=\frac{mv}{M}\\=\frac{0.012*410}{150}\\=0.0328m/s"